Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Combinatorics: Help Required

Expert replies
by Dblooos » Sun Jan 05, 2014 9:21 am
Team A plays two matches each with Team B and Team C. In any match, the probability of different outcome for Team A is given below:

Outcome Win / Loss / Draw


Probability 0.5 / 0.45 / 0.05


Points 2 / 0 / 1

Outcome of all the matches are independent of each other.

Q1: What is the probability of Team A getting at least 7 points in the contest? Assume Team B and Team C play 2 matches.
A. 0.025
B. 0.0875
C. 0.0625
D. 0.975
E None of these

Q2. What is the probability of Team C getting at least 4 points? Assume Team C and Team B play 2 matches.

A. 0.2025
B. 0.0625
C. 0.0425
D. 0.050
E. Can't be determined

My Doubts:

In question 1 the statement "Assume Team B and Team C play 2 matches" what information is this statement providing.

In question 2: "Assume Team C and Team B play 2 matches" is it not conflicting with the information in the us. Moreover the table only shows the probability of win or lose for Team A. So can I say it can't be determined?

My Solution Approach:

For Team A to have at least 7 points, out of total 4 matches that it plays with Team B and Team C. It should either Win 3 and draw 1 or Win all of them.

So the first case of winning all matches:

W X W X W X W = 0.5 X 0.5 X 0.5 X 0.5 = 0.0625

Second case, Winning 3 draw 1:

W X W X W X D = 0.5 X 0.5 X 0.5 X 0.05 = .00625 x 4 = .025 (as it is possible in following 4 ways)

W W W D
W W D W
W D W W
D W W W

So probability of getting at least 7 points = .025 + .0625 = 0.0875

Answer B;

Solution 2:

Can't determine as there is no information for the probability of outcome for the match between B & C. Is it correct?

Please advise. Thank you.
Join the discussion
Source: — Problem Solving |

by Mike@Magoosh » Mon Jan 06, 2014 4:47 pm
Dear Dblooos,
I'm happy to respond. :-)

I agree with your work, and agree that the answers should be
1) (B)
2) (E)
I believe this is what you meant.

Here's a blog with some additional challenging problems in probability:
https://magoosh.com/gmat/2014/gmat-advan ... -problems/

Enjoy!
Mike :-)
Magoosh GMAT Instructor
https://gmat.magoosh.com/
Join the discussion